OpenStudy (anonymous):

Solve: -2y < -6 or y + 3 ≤ 1

OpenStudy (anonymous):

--First: distribute everything. --Second: subtract 3y from each side. The 3y's on the left will cancel out. --Third: add 5 to each side. The 5's on the left side will cancel out. --Fourth: divide each side by 7. . 5(2y-1) = 3(y-4) 10y - 5 = 3y - 12 7y - 5 = -12 7y = -7 y = -1